OpenStudy (anonymous):

Meghan has a 15-year adjustable rate mortgage with a rate of 5.4% for the first 5 years. The monthly payment is $2,394.77. The amount of the mortgage is $295,000. What is the remaining balance after 5 years rounded to the nearest dollar? $253,392 $151,314 $221,673 $196,667 ---------------- I GOT $431,058.69 WHAT SHOULD DO FROM THERE ?

OpenStudy (anonymous):

Do I use this FVn = P[(1+c)n - 1]/c

OpenStudy (amistre64):

Balance time the (compounding set)^5; minus the payments times the geometric sum of the compunding setup

OpenStudy (amistre64):

i tend to be forgetting the scalar of 12 for some reason \[B_5=B_ok^{5*12}-P~\frac{1-k^{5*12}}{1-k}~:~k=1+\frac{.054}{12}\]

OpenStudy (anonymous):

So use the formula you used if I run into this problem again ?

OpenStudy (amistre64):

i tend to use that formula for all of it; i have to modify it at times tho

OpenStudy (anonymous):

whats the formula called ?

OpenStudy (amistre64):

dunno, i made it for myself ...

OpenStudy (anonymous):

lol can I get the original form of it to remember ?

OpenStudy (amistre64):

if we run thru what happens when you make payments on a balance:\[B_1=B_o(1+\frac r{12})-P\] \[B_2=B_1(1+\frac r{12})-P\]but replacing the value of B1 we get\[B_2=(B_0(1+\frac r{12})-P)(1+\frac r{12})-P\] for clean up lets say that 1+r/12 is k and continue \[B_2=(B_0(k)-P)k-P=B_0(k)^2-P(1+k)\] \[B_3=B_0(k)^3-P(1+k+k^2)\] \[B_4=B_0(k)^4-P(1+k+k^2+k^3)\] \[B_n=B_0(k)^n-P(1+k+k^2+k^3+...+k^{n-1})\]

OpenStudy (amistre64):

that tag along to the payments is a geometric sum with a formula of\[\frac{1-k^n}{1-k}\] therefore\[B_n=B_o(k)^n-P\frac{1-k^n}{1-k}\]

OpenStudy (amistre64):

this is the original form

OpenStudy (amistre64):

the n represent the number of periods, which tends to be adjusted as 12*number of years

OpenStudy (amistre64):

using this form i can calculate the payments need to zero out the balance after so many years by putting Bn = 0 and solving for P

OpenStudy (amistre64):

i can calculate annuities with it as well

OpenStudy (amistre64):

i can even use it to restructure thos ARM adjustable rate contrivances

OpenStudy (anonymous):

thank you soo much

OpenStudy (amistre64):

youre welcome, ive seen what the textbooks offer as formulas and they simply are not intuitive to me

OpenStudy (anonymous):

im confused @amistre64

OpenStudy (amistre64):

i use my own formula for this .... its one a created. But it works :)

OpenStudy (anonymous):

so wait... i got 221674, but thats not an optioon

OpenStudy (anonymous):

?

OpenStudy (amistre64):

221674 is an approximation ... 221673 is the closest one to it in the options

OpenStudy (anonymous):

oh, so its that one?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

the actual amount is: 221673.51

OpenStudy (anonymous):

kk ahhhhhhhhhhhhhhh